# LLMs can construct powerful representations and streamline sample-efficient supervised learning

### 12 Mar 2026 v1

@@ Abstract
As real-world datasets become increasingly complex and heterogeneous, supervised learning is often bottlenecked by input representation design. Modeling multimodal data for downstream tasks, such as time-series, free text, and structured records, often requires non-trivial domain-specific engineering. We propose an agentic pipeline to streamline this process. First, an LLM analyzes a small but diverse subset of text-serialized input examples in-context to synthesize a global rubric, which acts as a programmatic specification for extracting and organizing evidence. This rubric is then used to transform naive text-serializations of inputs into a more standardized format for downstream models. We also describe local rubrics, which are task-conditioned summaries generated by an LLM. Across 15 clinical tasks from the EHRSHOT benchmark, our rubric-based approaches significantly outperform traditional count-feature models, naive text-serialization-based LLM baselines, and a clinical foundation model, which is pretrained on orders of magnitude more data. Beyond performance, rubrics offer several advantages for operational healthcare settings such as being easy to audit, cost-effectiveness to deploy at scale, and they can be converted to tabular representations that unlock a swath of machine learning techniques.
Machine Learning, ICML
Figure 1: Performance averaged over all 15 clinical prediction tasks in the EHRSHOT benchmark with 6,739 patients (Wornow et al., 2023). Our rubric-style representations agentically constructed by LLMs outperform naive text-serialization-based LLM baseline in Hegselmann et al. (2025), as well as a clinical foundation model pretrained on 2.57M patients (CLMBR-T, (Wornow et al., 2023)), and a count feature-based gradient boosting machine (Count-GBM, (Ke et al., 2017; Wornow et al., 2023)).
@@ 1 Introduction
Supervised learning underpins a wide range of applications across domains. In medicine, deep neural networks achieve specialist-level performance in pneumonia detection and diabetic retinopathy screening (Rajpurkar et al., 2017; Gulshan et al., 2016). In finance, credit risk assessment models outperform legacy scorecards (Lessmann et al., 2015). In environmental science, supervised learning enables weather forecasting from radar observations (Ravuri et al., 2021).
1.1 Representational Challenges in Complex Domains
@@
A common thread across successful applications of supervised learning is the availability of input representations that can be easily processed by off-the-shelf models. Real-world datasets, however, are increasingly more complex and heterogeneous. They combine structured fields with unstructured text, time-stamped events, and different modalities such as images. In healthcare, clinical prediction may benefit from longitudinal labs and vitals, coded events (e.g., diagnoses and procedures), free-text notes, and medical images. In finance, stock-price forecasting and risk modeling may involve time series for trading prices and volumes, unstructured text (e.g., news and filings), and structured event records such as analyst rating changes.
@@
In domains where the input comprises a mix of heterogeneous and complex components, representation design requires substantial domain expertise and bespoke feature engineering, which may be unavailable. Even when expertise is available, resulting representations are not necessarily optimal: they may discard critical signal or bury it in noise, preventing the downstream model from harnessing it effectively. We show how large language models (LLM) can build agentic supervised learning pipelines that automate the design of powerful input representations and enable sample-efficient learning.
1.2 LLMs as a Representation Layer
LLMs offer a practical interface to heterogeneous data. Disparate components of the input can be serialized into text and directly processed with an LLM, bypassing domain-specific engineering. Recent work shows that LLMs can serve as effective predictors over complex inputs. Song et al. (2024) and Akhauri et al. (2025) serialize diverse system configurations and logs into unified text sequences to predict performance metrics and optimization outcomes. Hegselmann et al. (2025) serialize longitudinal electronic health records (EHR) into Markdown and train linear heads over their embeddings for clinical prediction tasks (see Figure 2, left, for an example of text-serialized EHR data). These works show LLMs’ potential to streamline supervised learning with complex datasets, but they treat text-serialization of input as fixed and leave the bulk of learning to the downstream model. In contrast, we take the text-serialized input as a starting point and show how LLMs can automate constructing better representations that dramatically improve downstream performance.
A complementary line of work is related to “data science (DS) agents”. Data Interpreter targets end-to-end benchmark-style problem solving, using decomposition, code generation, execution, and revision to complete data-analysis, machine-learning, and mathematical tasks (Hong et al., 2025). DS-Agent focuses on automating model-development workflows such as task understanding, model selection, and training (Guo et al., 2024). DeepAnalyze and DS-STAR push further toward autonomous data science over heterogeneous files, with an emphasis on multi-step data wrangling, open-ended querying, code execution, and report generation (Zhang et al., 2025b; Nam et al., 2025). This literature is closely related to our work in spirit, since it also uses LLMs as an interface to heterogeneous data. Our focus, however, is narrower and more controlled: rather than asking agents to plan and execute broad analyses, we study how LLMs can support representation design for supervised learning on complex data for specific downstream tasks. This lets us isolate the role of representation choice and demonstrate its effect as a first-order driver of downstream statistical performance in a controlled setting.
# Naive Text Serialization ## Patient Demographics
- Patient age: 78, FEMALE [...]
## Detailed Past Medical Visits ### Inpatient Visit (14 days to pred. time, current visit) #### Conditions
- Acute posthemorrhagic anemia
- pH measurement, venous: 7.25, 7.31, 7.31 [...] #### Medications
- furosemide 20 MG Oral Tablet
- pantoprazole 20 MG Delayed Release Oral Tablet [...] #### Procedures
- Chest x-ray
- Electrocardiogram report [...] ### Emergency Room Visit (87 days before prediction time) #### Conditions
- Benign essential hypertension
- Chest pain [...] #### Medications
- 2ML ondansetron 2MG/ML inject.
- nitroglycerin 0.4 MG [...] #### Procedures
- Ct angiography
- Comprehen metabolic panel[...]
@@ 1.2 LLMs as a Representation Layer
# Global Rubric Representation 3. Demographics - 55 | FEMALE | [...] 6. Recent Cardiac Symptoms (last 365 days) - Chest pain/angina: No - Dyspnea/shortness of breath: Yes (date unknown) [...] 12. Other Relevant Labs - Creatinine: 1.12 (2023-12-02) - eGFR: No data [...] 17. Known Risk Factors - Diabetes mellitus: No (A1c date unknown) - Hyperlipidemia: Yes - Family history of premature CAD: Unknown [...] 20. Non-cardiac Serious Illness That May Mimic or Alter MI Risk Interpretation - Active malignancy: No - Severe infection/sepsis in past 30 days: No - Major surgery in past 30 days: Yes — multiple inpatient procedures noted in December 2023 (e.g., CPT4/00520 on 2023-12-26 [...]
@@
Figure 2: Synthetic electronic health record (EHR) representation examples, focusing on the acute myocardial infarction (acute MI) prediction task. Left. Naive text-serialization adopted from Hegselmann et al. (2025). Middle. Local rubric representation which is a task-conditioned summary of the naive text-serialization. Right. Global rubric transformed version of the naive-text serialization.
1.3 A Frontier: Pre-trained Knowledge for Sample-efficiency
@@
Beyond data-processing convenience, LLMs possess pretraining knowledge that can enable effective regularization, which is key to sample-efficiency. Our work aligns with literature on injecting knowledge into statistical models: LMPriors uses language descriptions as task-specific priors (Choi et al., 2022), and TabLLM demonstrates effective few-shot tabular learning (Hegselmann et al., 2023). Similarly, LLM-Select and LLM-Lasso guide feature selection and regularization (Jeong et al., 2025; Zhang et al., 2025a), and Kim et al. (2025) use LLM-encoded task metadata to construct inductive biases.
@@
While these methods yield gains, they mostly use LLMs to augment traditional models working on clean datasets. In contrast, we focus on representation design: how complex inputs should be organized prior to downstream learning. In that sense, our work also aligns with recent literature on learning in the language space, such as GEPA by Agrawal et al. (2026).
1.4 Our Contribution: Rubric Representation Learning
We tackle the following question.
How can LLMs help construct powerful input representations in complex domains to streamline sample-efficient supervised learning for downstream tasks?
@@ 1.4 Our Contribution: Rubric Representation Learning
We propose rubric representation learning, where a team of LLM agents learn and implement rubrics, which are used to transform inputs into a standardized and information-rich format that can be easily and efficiently digested by downstream learners. We assume a naive text-serialization of the input is available, or it can be constructed straightforwardly (see Figure 2, left for an example). We develop two types of rubrics, which are given below and detailed in Section 2.
Global rubrics.
@@
We propose global rubrics, a novel input representation for complex datasets. A global rubric is a task-level specification that defines what information should be extracted from the input and how. It is generated by prompting an LLM with a diverse set of examples and asking it to produce a useful rubric, which is then used to transform the inputs (see Figure 2, right). We use global rubrics in two ways: (i) through vector embeddings of rubric-transformed text-serializations (ii) converting the rubric serialization into tabular features via an LLM (see Figure 3).
Local rubrics.
We prompt an LLM to produce a task-conditioned local summary with structured sections (see Figure 2, middle), similar to recent work on explainable clinical prediction models (Petridis et al., 2026). The summarization serves two key purposes: (i) it yields a less noisy representation that can be embedded more efficiently (ii) it injects the pretraining knowledge into the embeddings, delegating part of the learning to the upstream representation design step.
Advantages of global rubrics.
@@
Both rubrics achieve similar downstream performance and outperform the baselines. However, while all inputs share the same sections after the summarization step, local rubrics do not have the same level of standardization as global rubrics, which endows the latter with several practical desiderata lacking in the former.
@@
  • Auditable and improvable: Global rubrics are more amenable to inspections by domain experts, such as for analyzing subgroup bias risk, identifying missing or misleading evidence, and iterative refinement.
  • More operationally useful: Global rubric representations can be transformed to tabular features (Figure 3, Panel (F)), immediately enabling a suite of applications, such as interpretability methods and causal analyses. For instance, intervening on or controlling for variables are far simpler to do with tabular features than with free-form summaries.
@@
  • Cheaper to deploy at scale: Global rubric transformation at inference time can be automated (see Figure 3, Panels (E) and (F)), whereas summarization requires an LLM API call for each example. This can effectively make global rubrics “free” compared to local rubrics, which incur an ${\cal O}(N)$ cost in time and money. In contrast, global rubrics have ${\cal O}(1)$ monetary cost and can be applied orders of magnitude faster using deterministic scripts on CPU at inference time.
1.5 Evaluation in EHRSHOT Benchmark: Clinical Outcome Prediction
@@
We evaluate our methods on 15 clinical prediction tasks in the EHRSHOT benchmark, spanning four task groups: operational outcomes (3 tasks), assignment of new diagnoses (6 tasks), anticipating lab results (5 tasks), and prediction of chest X-ray findings (1 task) (Wornow et al., 2023). We compare against a gradient boosting machine (GBM) that uses count-based features (Count-GBM, (Ke et al., 2017)), a clinical foundation model pretrained on 2.57M patients (CLMBR-T, (Wornow et al., 2023)), zero-shot chain-of-thought prompting (CoT) with Qwen3-8B and GPT5-mini [^1] In compliance with the EHRSHOT license (Wornow et al., 2023), we used GPT5-Mini and GPT-5.2 via the HIPAA-compliant Microsoft Azure OpenAI Service. (Wei et al., 2022; Qwen3Team, 2025; OpenAI, 2025), and the LLM baseline in Hegselmann et al. (2025) which uses naive text-serializations of the patient EHRs. Our methods using rubric style representations outperform the baselines by a substantial margin on average. Relative to CLMBR-T, they achieve strong gains on new diagnosis and lab result tasks, and remain close for operational outcomes and chest X-ray findings tasks.
@@ 2.1 Global Rubrics
We introduce LLM-derived global rubrics, a domain-agnostic strategy for converting heterogeneous, weakly structured inputs into task-aligned representations that are more amenable to supervised learning. While our experiments focus on electronic health records (EHR) as the input, the procedure applies whenever inputs can be rendered as text.
@@
We describe the global rubric learning procedure for a single prediction task. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote labeled training data, where $x$ is a raw input and $y\in\{0,1\}$ is the task label. Let $s(\cdot)$ be some serialization procedure that maps an input to text and define $x^{\text{text}}=s(x)$. A rubric specifies a task-specific transformation
@@
where $x^{\text{rubric}}$ is a more structured representation of the same underlying input $x$, and it can be used with downstream predictors instead of $x^{\text{text}}$. We describe how $x^{\text{text}}$ and $x^{\text{rubric}}$ are used for downstream training in Section 4.
Global rubric synthesis.
Global rubric learning has two stages, which are shown in Figure 3, Panels (A) and (B). First, we select a small, label-balanced and diverse cohort from the training split. Second, an LLM inspects this cohort in-context and synthesizes a task-specific rubric by selecting and constructing predictive signals.
  • Step 1a) Diverse cohort selection (motivation):   Rubric synthesis is done through a single prompt to an LLM (GPT5-Mini) with a limited number of examples due to context length limitations (272k tokens). Since the in-context cohort drives rubric creation, cohort diversity is crucial for rubric to capture different modes of variation rather than reflecting a narrow set of training examples. Hence, we aim to build a cohort that (i) is label-balanced and (ii) spans a diverse region in training distribution.
  • Step 1b) Diverse cohort selection (procedure):   We first embed each text-serialized input $x_{i}^{\text{text}}$ into a vector space using a pretrained text embedding model (we use Qwen3-8B-Embedding (Zhang et al., 2025c)) and stratify by label,
    We perform $k$-means clustering independently within each stratum in the embedding space. Here, $k$ denotes the number of clusters per label stratum, so the cohort contains 2k2*k examples in total. Due to context windows limitations, we use $k=20$. From each cluster, we select the element that is closest to the center of the cluster (the medoid) into the cohort, yielding a compact cohort that covers diverse regions of the input space within both classes.
  • Step 2) Rubric synthesis:   Given the selected cohort, we ask an LLM to produce a task-specific rubric that (i) identifies discriminative, task-relevant signals, (ii) organizes them into coherent categories, and (iii) specifies how each signal should be extracted from a new input. The rubric standardizes how the input should be reorganized and transformed, but does not encode a prediction rule. The full rubric-synthesis prompt is provided in Appendix D.1, and two full-rubric examples can be found in Appendix E.1 and E.2.
@@
$\vdash$ Weight changes §5. ALERT FLAGS
$\vdash$ Resistant HTN markers
$\vdash$ End-organ damage [...]
@@
Figure 3: Agentic global-rubric pipeline for EHRSHOT tasks. (A) Build a label-balanced and diverse patient set via $k$-means. (B) Patient EHRs are fed to an LLM which is prompted to synthesize a task rubric. (C) The LLM outputs a systematic rubric $\mathcal{R}$ that defines how to transform any patient EHR from naive text ($x_{\text{text}}$) to textual rubric representation ($x_{\text{rubric}}$). (D) An LLM is asked to transform $x_{\text{text}}$ to $x_{\text{rubric}}$ for each patient. (E) An LLM is asked to write a script to automate the transformation step in Panel (D). (F) An LLM is asked to write a script to transform rubric representations $x_{\text{rubric}}$ into tabular features. Full prompts are provided in Appendix D.
Global rubric application.
Once a global rubric is learned, we apply it to naive text-serialization of each input, $x^{\text{text}}$, to produce $x^{\text{rubric}}$. We propose three different methods for the rubric transformation step, each with their own advantages.

### 20 May 2026 v3

@@ Abstract
As real-world datasets become more complex and heterogeneous, supervised learning is often bottlenecked by input representation design. Modeling multimodal data, such as time-series, free text, and structured records, often requires non-trivial domain expertise. We propose an agentic pipeline to streamline this process. First, an LLM analyzes a small but diverse subset of text-serialized input examples in-context to synthesize a global rubric, which acts as a programmatic specification for extracting and organizing evidence. This rubric is then used to transform naive text-serializations of inputs into a more standardized format for downstream models. We also describe local rubrics, which are task-conditioned interpretive summaries generated by an LLM. Across 15 clinical tasks from the EHRSHOT benchmark, our rubric approaches significantly outperform count-feature models, naive LLM baselines, and a clinical foundation model pretrained on orders of magnitude more data. Beyond performance, rubrics offer operational advantages such as being easy to audit, cost-effectiveness at scale, and facilitating tabular representations.
Figure 1: Performance (macro-averaged) over 15 clinical prediction tasks in the EHRSHOT benchmark [61], swept over training-set size $n$ (per task). Our rubric-style representations constructed by LLMs outperform the naive text-serialization baseline in Hegselmann et al. [15], as well as a clinical foundation model pretrained on 2.57M patient timelines (CLMBR-T, [61]), and a count feature-based gradient boosting machine (Count-GBM, [21, 61]).
@@ 1 Introduction
Supervised learning underpins a wide range of applications across domains. In medicine, deep neural networks achieve specialist-level performance in pneumonia detection and diabetic retinopathy screening [41, 12]. In finance, credit risk assessment models outperform legacy scorecards [26]. In environmental science, supervised learning enables weather forecasting from radar observations [42].
@@
The success of supervised learning depends on the availability of input representations that can be easily processed by off-the-shelf models. Real-world datasets, however, are increasingly more complex and heterogeneous. They combine structured fields with unstructured text, time-series, and images. In healthcare, clinical prediction may benefit from longitudinal labs, coded events (e.g., diagnoses), free-text notes, and medical images. In finance, stock-price forecasting and risk modeling may involve price and volume time-series, news reports, and structured events such as analyst ratings.
@@
In domains with complex data, representation design requires bespoke engineering and domain expertise. Even then, resulting representations are not necessarily optimal: they may discard critical signal or bury it in noise. We show how large language models (LLMs) can build agentic pipelines that automate designing powerful input representations and enable sample-efficient supervised learning.
LLMs offer a practical interface to heterogeneous data through text-serializations. Song et al. [50] and Akhauri et al. [3] serialize diverse configurations and logs into text sequences to predict system performance metrics. Hegselmann et al. [15] serialize longitudinal electronic health records (EHR) into Markdown and train linear heads over text-embeddings for clinical prediction tasks (see Figure˜2, left). Demirel et al. [8] use LLMs to predict daily user activities by combining multimodal time-series data from wearables after transforming them into textual descriptions.
These works show LLMs’ potential to streamline supervised learning with complex datasets, but they treat text-serialization of the input as fixed and leave the bulk of calibration to data for the downstream model. In contrast, we take the text-serialized input as a starting point and show how LLMs can automate constructing better representations to directly improve downstream performance.
@@
Naive Text Representation # Patient Demographics - Patient age: 78, Female […] # Detailed Past Medical Visits ## Inpatient Visit (14 days to pred. time, current visit) ### Conditions - Acute posthemorrhagic anemia - pH meas., venous: 7.25, 7.31 […] ### Medications - furosemide 20 MG Oral Tablet […] ### Procedures - Chest x-ray - Electrocardiogram report […] ## Emergency Room Visit (87 days before prediction time) ### Conditions - Benign essential hypertension - Chest pain […] Global Rubric Representation 3. Demographics - 55 | Female | […] 6. Recent Cardiac Symptoms (last 365 days) - Chest pain/angina: No - Dyspnea/shortness of breath: Yes (date unknown) […] 12. Other Relevant Labs - Creatinine: 1.12 (2023-12-02) - eGFR: No data […] 17. Known Risk Factors - Diabetes: No (A1c date unknown) - Family history of premature CAD: Unknown […] 20. Non-cardiac Serious Illness That May Mimic or Alter MI Risk - Active malignancy: No […]
@@
Figure 2: Synthetic electronic health record representation examples. Left. Naive text-serialization adopted from Hegselmann et al. [15]. Middle. Local rubric representation (task-conditioned summary of the naive text-serialization). Right. Global rubric transformed version of the naive-text serialization.
@@
Beyond modeling convenience, LLMs’ pretraining knowledge can enable effective regularization, which is key to sample-efficiency. Our work aligns with literature on injecting knowledge into statistical models: LMPriors uses language descriptions as task-specific priors [7], and TabLLM shows effective few-shot tabular learning [14]. LLM-Select and LLM-Lasso guide feature selection and regularization [17, 66], and Kim et al. [23] use task metadata to construct inductive biases.
@@
These methods mostly use LLMs to augment traditional models working on clean datasets. In contrast, we focus on representation design: how complex inputs should be organized prior to downstream learning. In that sense, our work also aligns with recent literature on learning in the language space, such as GEPA by Agrawal et al. [1]. An extended related work section is included in Appendix A.
@@
Our contribution.  We propose rubrics, which are used to process complex text inputs into a standardized and information-rich format that can be easily and efficiently digested by downstream learners. We assume naive text-serializations of inputs are available, or they can be constructed straightforwardly (see Figure˜2, left for an example). We develop two types of rubrics, which are given below and detailed in Section˜2.
@@
Global rubrics.  A global rubric is a task-level specification that defines what information should be extracted from the input and how. It is generated by prompting an LLM with a diverse set of examples (see Figure 3, Panels A–C).
Local rubrics.  We ask an LLM to produce a task-conditioned interpretive summary with structured sections (see Figure˜4, left), similar to recent work on explainable clinical prediction models [39].
@@
Advantages of global rubrics.  Both rubrics achieve similar downstream performance and outperform the baselines. However, local rubrics do not have the same level of standardization as global rubrics, which endows the latter with several practical desiderata lacking in the former.
@@
  • Auditable and improvable: Global rubrics are more amenable to inspections by domain experts, such as for analyzing subgroup bias risk and iterative refinement.
  • More operationally useful: Global rubric representations can be transformed into tabular features (Figure˜3, Panel F), immediately unlocking a suite of machine learning techniques.
@@
  • Cheaper to deploy: Global rubric transformation at inference time can be automated (see Figure˜3, Panels E and F), whereas summarization requires an LLM forward pass per example. This makes global rubrics “free” compared to local rubrics, which incur ${\cal O}(N)$ time and compute cost.
@@
We evaluate on 15 binary clinical prediction tasks from the EHRSHOT benchmark [61], spanning operational outcomes, new diagnoses, lab results, and chest X-ray findings. We compare against a gradient boosting machine with count-based features (Count-GBM, [21]), a clinical foundation model pretrained on 2.57M patients (CLMBR-T, [61]), zero-shot chain-of-thought prompting (CoT) with Qwen3-8B and GPT5-Mini [^1] We used GPT5-Mini and GPT-5.2 via the HIPAA-compliant Microsoft Azure OpenAI Service. [59, 40, 37], and the LLM baseline of Hegselmann et al. [15], which uses the naive EHR text-serializations we build on top of.
We preview the main findings here and provide detailed discussion in Section˜5. Rubrics substantially outperform baselines on average across sample sizes $n$, with gaps largest for small $n$. First, the LLM’s interpretation of evidence acts as a sample-efficiency lever. Stripping it from local rubrics costs noticeable performance at small $n$ and almost none at $n{=}\text{All}$, indicating that pretrained world knowledge supplies a prior the downstream classifier increasingly does without as data accumulate. Second, standardized global rubric templates are themselves strong representations even without an interpretive layer, beating all baselines across all sample sizes. Global rubrics trail local rubrics at small $n$, where the LLM-injected statistical prior in the language space matters most, but the gap closes by $n{=}\text{All}$ as labels accumulate.
@@ 2.1 Global rubrics
We introduce global rubrics for converting heterogeneous, weakly structured inputs into standardized, task-aligned representations. While we focus on electronic health records (EHR), the procedure applies wherever inputs can be text-serialized. Throughout, we use GPT5-Mini for natural-language steps (rubric synthesis and application, local-rubric generation) and GPT-5.2 for code-generation, as it produced more reliable scripts in our pilots (Figure˜3, Panels E, F).
@@
We describe the global rubric learning procedure for a single prediction task. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote labeled training data, where $x$ is a raw input and $y\in\{0,1\}$ is the task label. Let $s(\cdot)$ be some serialization procedure that maps an input $x$ into its textual representation, and define $x^{\text{text}}=s(x)$. A rubric specifies a task-specific transformation
@@
where $x^{\text{rubric}}$ is a more structured representation of the same underlying input $x$, and it can be used with downstream predictors instead of $x^{\text{text}}$. We describe downstream training in Section˜4.
@@
$\vdash$ Weight changes
@@
Figure 3: Agentic global-rubric pipeline for EHRSHOT tasks. Full prompts are in Appendix F.
Global rubric synthesis.
Global rubric learning has two stages, shown in Panels A and B of Figure˜3. First, we select a small, label-balanced and diverse cohort from the training split. Second, an LLM inspects this cohort in-context and synthesizes a task-specific rubric by defining predictive features and describing how to extract them.
  • Step 1) Diverse cohort selection:   Rubric synthesis is done through a single prompt to an LLM (GPT5-Mini). We first embed each text-serialized input $x_{i}^{\text{text}}$ into a vector space using a pretrained text-embedding model [69], and stratify by label,
    We perform $k$-means clustering within each stratum, where $k$ is the number of clusters per label-stratum, so the final cohort contains $2k$ examples. Due to context size limitations, we use $k=20$. Finally, we take the cluster medoids to obtain a diverse cohort (see Figure˜3, Panel A).
  • Step 2) Rubric synthesis:   Given the selected cohort, we ask an LLM (GPT5-Mini) to produce a task-specific rubric that (i) defines discriminative, task-relevant signals, and (ii) specifies how each signal should be extracted from a given input, $x^{\text{text}}$ (see Figure˜3, Panel B). The full prompt is provided in Appendix F.2, and two full-rubric examples can be found in Appendix G.
[…]